How much do wild mountain lions weigh? Adult wild mountain lions (18 months or older) captured and released for the first time in the San Andres Mountains gave the following weights (pounds):

66 103 125 122 60 64

Assume that the population of x values has an approximately normal distribution.

(a) Use a calculator with mean and sample standard deviation keys to find the sample mean weight x and sample standard deviation s. (Round your answers to one decimal place.)

x = lb
s = lb


(b) Find a 75% confidence interval for the population average weight μ of all adult mountain lions in the specified region. (Round your answers to one decimal place.)

lower limit     lb
upper limit     lb
Answer #1

Solution

\small \mathbf{x} (x-x) (x - x)
66 -24 576
103 13 169
125 35 1225
122 32 1024
60 -30 900
64 -26 676
\small \sum x=540 \small \small \sum \mathbf{\left (x-\bar x \right )}^{2}=4570

\small {\color{Blue} \mathbf{(a)}}

\small \mathbf{Sample\;Mean\;(\bar X)=\frac{\sum X}{n}=\frac{540}{6} =\mathbf{{\color{Red} 90.0}}\;lb}

\small \small \mathbf{Standard\;deviation\;(s)}=\sqrt{\frac{\sum \left (x-\bar x \right )^{2}}{n-1}}=\sqrt{\frac{4570}{6-1}}=\mathbf{{\color{Red} 30.2}\;lb}

------------------------------------------------------------------------------------------------------------

\small \mathbf{{\color{Blue} (b)}\;Confidence\;level\;(C)=0.75}

\small \alpha =1-C=1-0.75=\mathbf{0.25}

\small \small \mathbf{degrees\;of\;freedom\;(df)}=n-1=6-1=\mathbf{{\color{Red} }5}

Refer t-table or use excel function "=T.INV.2T(0.25,5)" to find the critical value of t.

\small \mathbf{\therefore t_{\frac{\alpha }{2}}=2.015 }

formula : confidence interval for the population mean. 11

{ } - 1+=

\small =90.0 \pm 2.015*\left \{ \frac{30.2}{\sqrt{6}} \right \}

\small =90.0 \pm 24.9

\small \mathbf{\mathbf{lower\;limit}=90.0-24.9={\color{Red} 65.1}\;lb}

\small \mathbf{\mathbf{upper\;limit}=90.0+24.9={\color{Red} 114.9}\;lb}